OpenStudy (anonymous):

Can Any one help me with this question? I have a exam tomorrow.. Q) A rectangular sheet of metal measuring 120cm by 75 cm has equal squares of side x cut from each of the corners. The Remaining flaps are then folded to form an open box. Find the value of x such the volume is a maxim. Thanks in Advance !!

OpenStudy (anonymous):

 volume of the box will be = (120-2x)*(75-2x)*x= =9000x-240x^2-150x^2+4x^3=4x^3-390x^2+9000x to find the maximum value of this function, find 1º derivative and look for values of x which make it 0. If need more help tell me, :)